The question is to solve algebraically:
x-3y-3z=0
2x+5y-5z=1
-x+5y-6z=-9
After I solved for the first 2 equations i got 11y-11Z=1 then I solved for the second and third equation and got 15y-129z=-8 then i combined them and got 26y+140z=-7 but the objective is to find out what x,y,z is and I was not successful in that

Answer by Alan3354(69443)

x-3y-3z=0
2x+5y-5z=1
-x+5y-6z=-9
After I solved for the first 2 equations i got 11y-11Z=1 then I solved for the second and third equation and got 15y-129z=-8 then i combined them and got 26y+140z=-7 but the objective is to find out what x,y,z is and I was not successful in that
==================
11y-11Z=1
15y-129z=-8
I don't know how you got those.
--------------
x-3y-3z=0
2x+5y-5z=1
-x+5y-6z=-9
---
Add the 1st and 3rd
2y - 9z = -9 Eqn A
----
2x-6y-6z=0 1st eqn times 2
2x+5y-5z=1 2nd eqn
----------------- Subtract
-11y - z = -1 Eqn B
99y + 9z = 9 Eqn B times 9
2y - 9z = -9 Eqn A
----------------------------- Add
101y = 0
y = 0
-----------------
2y - 9z = -9 Eqn A
z = 1
---------
Sub for y & z in any eqn to find x